2x + y = 7

x - 2y = 1


A. Write the equation in matrix form.

B. Determine the inverse of the matrix

C. Hence solve the equations.

D. x and y are matrices


"X=\begin{bmatrix}
 1 & 5 \\
 3 & 7
\end{bmatrix} 
 

 Y=\begin{bmatrix}
 3 & 4 \\
 2 & 1
\end{bmatrix}"

Evaluate X2 + Y


1
Expert's answer
2022-03-30T17:10:32-0400

"A:\\\\\\left[ \\begin{matrix}\t2&\t\t1\\\\\t1&\t\t-2\\\\\\end{matrix} \\right] \\left[ \\begin{array}{c}\tx\\\\\ty\\\\\\end{array} \\right] =\\left[ \\begin{array}{c}\t7\\\\\t1\\\\\\end{array} \\right] \\\\B:\\\\\\left[ \\begin{matrix}\t2&\t\t1\\\\\t1&\t\t-2\\\\\\end{matrix} \\right] =\\frac{1}{2\\cdot \\left( -2 \\right) -1\\cdot 1}\\left[ \\begin{matrix}\t-2&\t\t-1\\\\\t-1&\t\t2\\\\\\end{matrix} \\right] =\\left[ \\begin{matrix}\t0.4&\t\t0.2\\\\\t0.2&\t\t-0.4\\\\\\end{matrix} \\right] \\\\C:\\\\\\left[ \\begin{array}{c}\tx\\\\\ty\\\\\\end{array} \\right] =\\left[ \\begin{matrix}\t0.4&\t\t0.2\\\\\t0.2&\t\t-0.4\\\\\\end{matrix} \\right] \\left[ \\begin{array}{c}\t7\\\\\t1\\\\\\end{array} \\right] =\\left[ \\begin{array}{c}\t3\\\\\t1\\\\\\end{array} \\right] \\\\D:\\\\X^2+Y=\\left[ \\begin{matrix}\t1&\t\t5\\\\\t3&\t\t7\\\\\\end{matrix} \\right] \\left[ \\begin{matrix}\t1&\t\t5\\\\\t3&\t\t7\\\\\\end{matrix} \\right] +\\left[ \\begin{matrix}\t3&\t\t4\\\\\t2&\t\t1\\\\\\end{matrix} \\right] =\\left[ \\begin{matrix}\t16&\t\t40\\\\\t24&\t\t64\\\\\\end{matrix} \\right] +\\left[ \\begin{matrix}\t3&\t\t4\\\\\t2&\t\t1\\\\\\end{matrix} \\right] =\\left[ \\begin{matrix}\t19&\t\t44\\\\\t26&\t\t65\\\\\\end{matrix} \\right]"